OpenStudy (anonymous):

(1+sinx)/cosx + (cosx/(1+sinx)=2secx

OpenStudy (anonymous):

just add the fractions together and simplify (use cos^2 + sin^2 = 1)

OpenStudy (anonymous):

im still confused as to how i add these together.

OpenStudy (anonymous):

Just take the denominator as cosx(1+sinx) and proceed as usual....

OpenStudy (anonymous):

oh, so i can leave it like that without distributing it?

OpenStudy (anonymous):

so far i get: (1+sin+cos)/(cos)(1+sin)

OpenStudy (anonymous):

Your numerator is not correct. Just do it as if they were normal fractions.

OpenStudy (asnaseer):

use this fact:\[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\]

OpenStudy (anonymous):

i get 1+2sin+sin^2+cos^2. does this look better for the numerator?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

As I said in my first post, now simplify (and use cos^2 + sin^2 = 1)

OpenStudy (anonymous):

(1+2sin+sin^2+cos^2)/(cos)(1+sin) so far

OpenStudy (anonymous):

now i get (2+2sin)/(cos)(1+sin)

OpenStudy (anonymous):

There is a fairly obvious simplification there....

OpenStudy (anonymous):

You are almost done....